### v3

@@ 1 Introduction
Volatility, defined as the standard deviation or variance of asset returns, evolves stochastically over time and plays a central role in financial risk management. Accurate volatility forecasting is essential for evaluating tail risks such as value-at-risk (VaR) and expected shortfall (ES). Traditionally, two major classes of time-series models have been used for modeling time-varying volatility: the generalized autoregressive conditional heteroskedasticity (GARCH) family (Engle, 1982; Bollerslev, 1986), and the stochastic volatility (SV) model (Taylor, 1986). These models successfully capture key features of financial volatility, including volatility clustering and persistence, and have been extended to account for leverage effects—i.e., the negative correlation between current returns and future volatility. A prominent example is the exponential GARCH (EGARCH) model by Nelson (1991).
@@ 2.1 Stochastic Volatility Model
where $p_{t}$ is the closing price on day $t$. The return variance is modeled as
$\mathrm{E}[y_{t}^{2}]=\sigma_{t}^{2}=\exp(h_{t}),$
where $h_{t}$ denotes the log volatility, treated as a latent process.
@@ 2.2 Realized Volatility
[…]
Assume that $n$ intraday returns are observed on day $t$, denoted by $\{r_{t},r_{t+1/n},\ldots,r_{t+(n-1)/n}\}$. Then, the realized volatility is defined as
$RV_{t}=\sum_{i=0}^{n-1}r_{t+i/n}^{2}.$ (8)
Under ideal conditions (e.g., no microstructure noise or missing observations), $RV_{t}$ converges to $IV_{t}$ as $n\to\infty$.
@@ 2.3 Realized Stochastic Volatility Model
The parameter $\xi$ adjusts for systematic bias in $\log RV_{t}$. When $\xi=0$, RV is an unbiased estimate of the integrated volatility. In practice, however, $RV_{t}$ may be biased downward due to non-trading hours and upward due to microstructure noise. Hansen and Lunde (2006) show that the net bias can be either positive or negative. Therefore, the sign of $\xi$ reflects the relative magnitude of these two opposing effects.
@@ 2.4 RSV Model with Student’s $t$ Distribution
[…]
@@ 2.5 RSV Model with GH Skew-$t$ Distribution
The denominator in (25) standardizes $\epsilon_{t}$ so that its conditional variance equals $\exp(h_{t})$. When $\beta=0$, the model reduces to the RSV-T specification based on the standard Student’s $t$ distribution.
In this formulation, the skewness and tail thickness of $\epsilon_{t}$ are jointly governed by the parameters $\beta$ and $\nu$. The parameter $\beta$ introduces asymmetry, while $\nu$ controls tail heaviness. As $\nu\to\infty$, or if $\lambda_{t}=1$ for all $t$, the distribution converges to the Gaussian case regardless of $\beta$.
@@
To illustrate the influence of $\beta$ and $\nu$, Figure 1 plots the simulated densities of $\epsilon_{t}$ under various parameter values. Panel (i) shows that increasing negative values of $\beta$ yield more pronounced left-skewness and heavier tails. In contrast, Panel (ii) indicates that as $\nu$ increases, the distribution becomes more symmetric and approaches normality, exhibiting thinner tails.
Figure 1: Simulated density of the standardized GH skew-$t$ distribution in equation (25). (i) Varying $\beta=0,-1,-2$ with fixed $\nu=10$. (ii) Varying $\nu=24,16,8$ with fixed $\beta=-1$.
@@ 2.6 RSV Model with Azzalini Skew-$t$ Distribution
Figure 2: Simulated density of the standardized Azzalini skew-$t$ distribution in equation (30). (i) $\delta=0,-0.6,-0.9$ with $\nu=10$ fixed. (ii) $\nu=15,10,5$ with $\delta=-0.9$ fixed.
@@ 2.7 RSV Model with Fernández–Steel Skew-$t$ Distribution
[…]
@@ 3.1 RSV-AZ-ST Model
$\frac{\delta+1}{2}$ $\sim\mathcal{B}(a_{\delta 0},b_{\delta 0}),\quad\nu\sim\mathcal{G}(n_{\nu 0},S_{\nu 0}),$ (50)
@@ 3.2 RSV-FS-ST Model
$\gamma\sim\mathcal{G}(n_{\gamma 0},S_{\gamma 0}),\quad\nu\sim\mathcal{G}(n_{\nu 0},S_{\nu 0}).$ (51)
@@ 3.3 One-day-ahead Forecast
where $x$ and $f$ represent a volatility proxy and a volatility forecast, respectively. The QLIKE loss function is robust, as defined by Patton (2011), meaning that it provides consistent rankings regardless of whether the ranking is based on true volatility or a conditionally unbiased volatility proxy. Furthermore, Patton and Sheppard (2009) demonstrated its superior power compared to the mean squared error (MSE), another robust loss function, in the predictive precision test proposed by Diebold and Mariano (2002).
@@
where $\hat{\Omega}_{n_{f}}=n_{f}^{-1}\sum_{t=n}^{n+n_{f}-1}Z_{t+1}Z_{t+1}^{\prime}$ consistently estimates the variance of $Z_{t+1}$. Under the null, $T_{n_{f}}$ follows a chi-square distribution with $q$ degrees of freedom, $\chi_{q}^{2}$. At the significance level $p$, the null hypothesis is rejected if $T_{n_{f}}>\chi_{q,1-p}^{2}$, where $\chi_{q,1-p}^{2}$ is the $(1-p)$ quantile of $\chi_{q}^{2}$.
@@ 4.1 Data and Descriptive Statistics
Table 1: Descriptive statistics of daily returns for the DJIA and N225
@@
Figure 5 shows the time series and histograms of the 5-minjte RVs and their logarithmic transformations for both indices. Both series exhibit strong temporal clustering, high persistence, and occasional sharp spikes. [^5] For the DJIA, the largest RV spike occurred on August 24, 2015, amid heightened market turbulence (see The New York Times: https://www.nytimes.com/2015/08/25/business/dealbook/daily-stock-market-activity.html). For the N225, the most notable spike on March 15, 2011, reflects the aftermath of the Great East Japan Earthquake on March 11, 2011. These spikes lead to positively skewed distributions in the log-RVs.
@@
Table 2 summarizes the descriptive statistics of the log-RVs. The Ljung–Box test strongly rejects the null of no autocorrelation, affirming the presence of volatility clustering in both markets. The distributions are positively skewed and leptokurtic. Consequently, the JB test rejects the normality of the log-RVs. This finding is inconsistent with the normality assumption for the error term $u_{t}$ in equation (16). While we maintain this assumption for tractability, allowing for non-normality in future model extensions would be a valuable research direction.
Table 2: Descriptive statistics of the logarithms of 5-minute RVs for the DJIA and N225
@@
  • Notes: See Table 1 for additional details.

### v4

@@ 1 Introduction
Volatility, defined as the conditional standard deviation or variance of asset returns, evolves stochastically over time and plays a central role in financial risk management. Accurate volatility forecasting is essential for evaluating tail risks such as value-at-risk (VaR) and expected shortfall (ES). Traditionally, two major classes of time-series models have been used for modeling time-varying volatility: the generalized autoregressive conditional heteroskedasticity (GARCH) family (Engle, 1982; Bollerslev, 1986), and the stochastic volatility (SV) model (Taylor, 1986). These models successfully capture key features of financial volatility, including volatility clustering and persistence, and have been extended to account for leverage effects—i.e., the negative correlation between current returns and future volatility. A prominent example is the exponential GARCH (EGARCH) model by Nelson (1991).
@@
Score-driven (SD) models, introduced by Creal et al. (2013) and Harvey (2013), have also gained popularity as a flexible alternative to SV and GARCH. [^1] An extensive and regularly updated list of SD models proposed in the literature is available at https://www.gasmodel.com/. Catania and Grassi (2022) apply the GH skew-$t$ distribution within an SD–GARCH setting, while Catania and Proietti (2020) develop a joint SD model for returns and realized volatility using a time-varying bivariate $t$ specification. These approaches are conceptually related to ours, but their dynamic updating rules and likelihood evaluations are substantially more complex. We therefore focus on SV/RSV-type models, while acknowledging SD-based specifications as an important complementary direction in the broader volatility-modeling literature.
@@ 2.1 Stochastic Volatility Model
where $p_{t}$ is the closing price on day $t$. The conditional return variance is modeled as
$\mathrm{E}[y_{t}^{2}\mid\mathcal{I}_{t-1}]=\sigma_{t}^{2}=\exp(h_{t}),$
where $\mathcal{I}_{t-1}$ denotes the information set available at time t1t-1 and $h_{t}$ denotes the conditional log volatility, treated as a latent process.
@@ 2.2 Realized Volatility
$IV_{t}=\int_{t-1}^{t}\sigma^{2}(s)ds.$ (7)
Assume that $m$ intraday returns are observed on day $t$, denoted by $\{r_{t-1+1/m},r_{t-1+2/m},\ldots,r_{t}\}$. Then, the realized volatility is defined as
$RV_{t}=\sum_{i=1}^{m}r_{t-1+i/m}^{2}.$ (8)
Under ideal conditions (e.g., no microstructure noise or missing observations), $RV_{t}$ converges to $IV_{t}$ as $m\to\infty$.
@@ 2.3 Realized Stochastic Volatility Model
The parameter $\xi$ adjusts for systematic bias in $x_{t}=\log RV_{t}$. When $\xi=0$, $x_{t}$ is an unbiased estimate of $h_{t}$ . In practice, however, $RV_{t}$ may be biased downward due to non-trading hours and upward due to microstructure noise. Hansen and Lunde (2006) show that the net bias can be either positive or negative. Therefore, the sign of $\xi$ reflects the relative magnitude of these two opposing effects.
@@ 2.4 RSV Model with Student’s $t$ Distribution
$\lambda_{t}$ $\overset{\text{i.i.d.}}{\sim}\mathcal{IG}\left(\frac{\nu}{2},\frac{\nu}{2}\right),\quad\mu_{\lambda}=\mathrm{E}[\lambda_{t}]=\frac{\nu}{\nu-2},\quad\nu>2.$ (22)
@@ 2.5 RSV Model with GH Skew-$t$ Distribution
A distinctive feature of the GH skew-$t$ distribution, originally emphasized by Aas and Haff (2006), is its asymmetric tail behavior: one tail exhibits polynomial (power-law) decay, while the opposite tail decays exponentially, depending on the sign of the skewness parameter $\beta$. This feature makes the GH skew-$t$ distribution particularly flexible in capturing return asymmetry and tail risk in financial applications.
The denominator in (25) standardizes $\epsilon_{t}$ so that its conditional variance equals $\exp(h_{t})$. In the present formulation, the skewness and tail behavior of $\epsilon_{t}$ are jointly governed by $\beta$ and $\nu$. The parameter $\beta$ introduces asymmetry and determines which tail becomes heavier, while $\nu$ controls the degree of polynomial tail thickness. When $\beta=0$, the model reduces to the RSV-T specification in . As $\nu\to\infty$, or if $\lambda_{t}=1$ for all $t$, the distribution converges to the Gaussian case regardless of $\beta$.
@@
To illustrate the influence of $\beta$ and $\nu$, Figure 1 plots the simulated densities of $\epsilon_{t}$ under various parameter values. Panel (i) shows that increasing negative values of $\beta$ yield more pronounced left-skewness and heavier polynomial tails. In contrast, Panel (ii) indicates that as $\nu$ increases, the distribution becomes more symmetric and approaches normality, exhibiting thinner tails.
Figure 1: Density of the standardized GH skew-$t$ distribution in equation (25). (i) Varying $\beta=0,-1,-2$ with fixed $\nu=10$. (ii) Varying $\nu=24,16,8$ with fixed $\beta=-1$.
@@ 2.6 RSV Model with Azzalini Skew-$t$ Distribution
Figure 2: Density of the standardized Azzalini skew-$t$ distribution in equation (30). (i) $\delta=0,-0.6,-0.9$ with $\nu=10$ fixed. (ii) $\nu=15,10,5$ with $\delta=-0.9$ fixed.
@@ 2.7 RSV Model with Fernández–Steel Skew-$t$ Distribution
$M_{1}$ $=2\int_{0}^{\infty}\tilde{w}f_{T}(\tilde{w}\mid\nu)d\tilde{w}=\frac{2c_{\nu}\nu}{\nu-1},$ (42)
$M_{2}$ $=2\int_{0}^{\infty}\tilde{w}^{2}f_{T}(\tilde{w}\mid\nu)d\tilde{w}=\frac{\nu}{\nu-2}.$ (43)
@@
Following Trottier and Ardia (2016), higher-order moments for the FS skew-$t$ distribution and their standardized forms are derived utilizing the absolute moments of the underlying Student’s $t$ distribution. For the absolute moments, we refer to the unified results provided by Kirkby et al. (2025). The $k$-th absolute moment, denoted by $M_{k}=2\int_{0}^{\infty}\tilde{w}^{k}f_{T}(\tilde{w}\mid\nu)d\tilde{w}$, is generally given by:
$M_{k}=\nu^{k/2}\frac{\Gamma\left(\frac{k+1}{2}\right)\Gamma\left(\frac{\nu-k}{2}\right)}{\sqrt{\pi}\,\Gamma\left(\frac{\nu}{2}\right)},\quad k<\nu.$ (45)
Specifically, for $k=3$ and $k=4$, we obtain:
$M_{3}$ $=\frac{\nu^{3/2}\,\Gamma\left(\frac{\nu-3}{2}\right)}{\sqrt{\pi}\,\Gamma\left(\frac{\nu}{2}\right)}\quad(\nu>3),$ (46)
$M_{4}$ $=\frac{3\nu^{2}}{(\nu-2)(\nu-4)}\quad(\nu>4).$ (47)
Using the general raw-moment formula provided in Trottier and Ardia (2016):
$\mu_{r}^{\prime}=E[w^{r}]=M_{r}\frac{\gamma^{r+1}+(-1)^{r}\gamma^{-(r+1)}}{\gamma+\gamma^{-1}},\quad r=1,2,3,4,$
the centered third and fourth moments are defined as:
$\mu_{3}=\mu_{3}^{\prime}-3\mu_{*}\mu_{2}^{\prime}+2\mu_{*}^{3},\qquad\mu_{4}=\mu_{4}^{\prime}-4\mu_{*}\mu_{3}^{\prime}+6\mu_{*}^{2}\mu_{2}^{\prime}-3\mu_{*}^{4}.$
Consequently, the skewness and excess kurtosis of the standardized variable $w_{*}=(w-\mu_{*})/\sigma_{*}$ are given by:
$\mathrm{Skew}(w_{*})$ $=\frac{\mu_{3}}{\sigma_{*}^{3}},$ (48)
$\mathrm{Kurt}(w_{*})$ $=\frac{\mu_{4}}{\sigma_{*}^{4}}-3.$ (49)
These expressions implicitly depend on $(\gamma,\nu)$ through the moments $M_{k}$ and the parameter $\gamma$, confirming that $\gamma$ governs asymmetry while $\nu$ determines tail thickness.
@@
2.8 EGARCH Model
GARCH models are widely used to describe time-varying volatility. Among numerous variants, we consider an EGARCH-type model. While the original EGARCH specification of Nelson (1991) involves the term $\mathrm{E}[|\epsilon_{t}|]$, this expectation does not admit a convenient closed-form expression under standardized $t$ innovations. Following Hansen and Huang (2016), we therefore estimate the model using the quadratic term $(\epsilon_{t}^{2}-1)$ instead, which yields
$y_{t}$ $=\epsilon_{t}\exp\!\left(\frac{h_{t}}{2}\right),\quad\epsilon_{t}\sim\mathcal{N}(0,1),\quad t=1,\ldots,n,$ (52)
$h_{t+1}$ $=\omega+\varphi(h_{t}-\omega)+\tau\epsilon_{t}+\gamma(\epsilon_{t}^{2}-1),\quad|\varphi|<1,\quad t=1,\ldots,n-1,$ (53)
with the initial condition $h_{1}=\omega$. The parameter $\tau$ captures the asymmetric response of conditional volatility to return shocks, and when $\tau<0$, negative shocks increase volatility more than positive ones, reflecting the leverage effect. For expositional clarity, we present the model above under the Gaussian assumption for $\epsilon_{t}$. In Section 4, we relax this assumption and consider a standardized Student-t distribution with degrees-of-freedom parameter $\nu$, allowing for heavy-tailed return behavior.
Remark.
Nelson (1991) noted that estimating EGARCH with Student-t innovations can suffer from stability issues (p. 365), and suggested the generalized error distribution (GED) as a more robust alternative. While Nelson (1991) originally proposed a quasi-maximum likelihood (QML) estimation procedure, in this study we estimate the model parameters using a Bayesian approach. We acknowledge the importance of the EGARCH-GED specification; however, given the additional computational complexity of Bayesian estimation and the main goal of this study to compare realized volatility-based stochastic volatility models, we leave the estimation of EGARCH-GED for future work.
2.9 REGARCH Model
Hansen et al. (2012) and Hansen and Huang (2016) extend the conventional GARCH and EGARCH models to the RGARCH and REGARCH frameworks, respectively, by incorporating realized measures into the volatility dynamics. Since the REGARCH model allows a more flexible characterization of the joint dynamics between returns and volatility than the RGARCH model, we adopt the REGARCH specification, which can be written as
$y_{t}$ $=\epsilon_{t}\exp\!\left(\frac{h_{t}}{2}\right),\quad\epsilon_{t}\sim\mathcal{N}(0,1),\quad t=1,\ldots,n,$ (54)
$h_{t+1}$ $=\omega+\varphi(h_{t}-\omega)+\tau_{1}\epsilon_{t}+\tau_{2}(\epsilon_{t}^{2}-1)+\gamma\upsilon_{t},\quad|\varphi|<1,\quad t=1,\ldots,n-1,$ (55)
$x_{t}$ $=\zeta+h_{t}+\delta_{1}\epsilon_{t}+\delta_{2}(\epsilon_{t}^{2}-1)+\upsilon_{t},\quad\upsilon_{t}\sim\mathcal{N}(0,\sigma_{\upsilon}^{2}),\quad t=1,\ldots,n,$ (56)
with the initial condition $h_{1}=\omega$. Similarly to the EGARCH model, in Section 4 we also consider a standardized Student-t distribution for $\epsilon_{t}$ in the REGARCH model, with degrees-of-freedom parameter $\nu$, allowing for heavy-tailed return behavior.
The latent log-volatility process $h_{t}$ and the realized measure $x_{t}$ each include a quadratic form of the leverage function, given by $\tau_{1}\epsilon_{t}+\tau_{2}(\epsilon_{t}^{2}-1)$ and $\delta_{1}\epsilon_{t}+\delta_{2}(\epsilon_{t}^{2}-1)$, respectively. Following Hansen and Huang (2016), one can introduce an additional scaling parameter $\psi$ as
$x_{t}=\zeta+\psi h_{t}+\delta_{1}\epsilon_{t}+\delta_{2}(\epsilon_{t}^{2}-1)+\upsilon_{t},$
but since the estimate of $\psi$ is typically close to unity, we fix $\psi=1$ for brevity. The parameter $\zeta$ captures the bias in $\log RV_{t}$; $\zeta=0$ implies an unbiased realized measure.
@@ 3.1 RSV-AZ-ST Model
$\frac{\delta+1}{2}$ $\sim\mathcal{B}(a_{\delta 0},b_{\delta 0}),\quad\nu\sim\mathcal{G}(n_{\nu 0},S_{\nu 0})\mathbbm{1}\{\nu>4\},$ (60)
@@ 3.2 RSV-FS-ST Model
$\gamma\sim\mathcal{G}(n_{\gamma 0},S_{\gamma 0}),\quad\nu\sim\mathcal{G}(n_{\nu 0},S_{\nu 0})\mathbbm{1}\{\nu>4\}.$ (61)
@@ 3.3 One-day-ahead Forecast
where $x$ and $f$ represent a volatility proxy and a volatility forecast, respectively. The QLIKE loss function is robust, as defined by Patton (2011), meaning that it provides consistent rankings regardless of whether the ranking is based on true volatility or a conditionally unbiased volatility proxy. Furthermore, Patton and Sheppard (2009) demonstrated its superior power compared to the mean squared error (MSE), another robust loss function, in the predictive accuracy test proposed by Diebold and Mariano (1995).
@@
where $\hat{\Omega}_{n_{f}}=n_{f}^{-1}\sum_{t=n}^{n+n_{f}-1}Z_{t+1}Z_{t+1}^{\prime}$ consistently estimates the variance of $Z_{t+1}$. Under the null, $T_{n_{f}}$ follows a chi-square distribution with $q$ degrees of freedom, $\chi_{q}^{2}$. At the significance level $p$, the null hypothesis is rejected if $T_{n_{f}}>\chi_{q,1-p}^{2}$, where $\chi_{q,1-p}^{2}$ is the (1p)(1-p) quantile of $\chi_{q}^{2}$.
@@
3.4 EGARCH and REGARCH Models
Following Mitsui and Watanabe (2003), we use the Metropolis–Hastings (MH) algorithm to sample the parameters of the EGARCH and REGARCH models from the posterior distribution. [^4] For other Bayesian approaches to RGARCH-type models and their applications to volatility and tail risk forecasting, see Chen et al. (2023). First, we obtain the mode of the posterior density, and then set the proposal density to a multivariate normal distribution centered at the posterior mode, with a covariance matrix proportional to the inverse Hessian of the log posterior density evaluated at the mode, scaled by a factor of 1.2. The scaling factor is introduced to ensure sufficient exploration of the tails of the posterior distribution. Asai (2006) shows that this method is efficient for sampling the parameters of GARCH-type models from the posterior distribution.
The prior distributions for the EGARCH model are specified as follows:
$\omega\sim\mathcal{N}(0,100),\quad\frac{\varphi+1}{2}\sim\mathcal{B}(1,1),\quad-\tau\sim\mathcal{G}(0.1,0.1),\quad\gamma\sim\mathcal{G}(0.1,0.1),\quad\nu\sim\mathcal{G}(5,0.5)\mathbb{I}(\nu>4),$ (73)
For the additional parameters of the REGARCH model, the prior distributions are given by
$\tau_{1}\sim\mathcal{G}(0.1,0.1),\quad\tau_{2}\sim\mathcal{G}(0.1,0.1),\quad\zeta\sim\mathcal{N}(0,10),\quad-\delta_{1}\sim\mathcal{G}(0.1,0.1),\quad\delta_{2}\sim\mathcal{G}(0.1,0.1),\quad\sigma_{\upsilon}^{2}\sim\mathcal{IG}(2.5,0.5).$
The parameters $\tau$, $\tau_{1}$, $\tau_{2}$, $\gamma$, $\delta_{1}$, and $\delta_{2}$ are, in principle, unrestricted in sign. However, empirical evidence typically suggests $\tau<0$ for EGARCH, and $\tau_{1}<0$, $\tau_{2}>0$, $\gamma>0$, $\delta_{1}<0$, and $\delta_{2}>0$ for REGARCH models. Accordingly, we employ Gamma priors with appropriate sign restrictions to confine these parameters to economically meaningful regions.
Based on the posterior draws, one-day-ahead forecasts of volatility, VaR, and ES are obtained. This approach allows us to jointly sample both parameters and the implied conditional volatilities from the posterior distribution, thereby explicitly incorporating parameter uncertainty into volatility and return forecasts.
Importantly, this feature addresses concerns raised in the literature that ignoring parameter uncertainty in GARCH-type models may lead to unfair comparisons with stochastic volatility models; see, for example, Ardia et al. (2018), who document that accounting for parameter uncertainty improves volatility and risk forecasts in GARCH frameworks. By adopting a fully Bayesian estimation strategy, our framework ensures that parameter uncertainty is treated consistently across models, providing a fair basis for comparison and yielding coherent predictive distributions for one-step-ahead volatility and return forecasts.
@@ 4.1 Data and Descriptive Statistics
Table 1: Descriptive statistics of daily returns for the DJIA and N225.
@@
Figure 5 shows the time series and histograms of the 5-minute RVs and their logarithmic transformations for both indices. Both series exhibit strong temporal clustering, high persistence, and occasional sharp spikes. [^7] For the DJIA, the largest RV spike occurred on August 24, 2015, amid heightened market turbulence (see The New York Times: https://www.nytimes.com/2015/08/25/business/dealbook/daily-stock-market-activity.html). For the N225, the most notable spike on March 15, 2011, reflects the aftermath of the Great East Japan Earthquake on March 11, 2011. These spikes lead to positively skewed distributions in the log-RVs.
@@
Table 2 summarizes the descriptive statistics of the log-RVs. The Ljung–Box test strongly rejects the null of no autocorrelation, confirming the presence of volatility clustering in both markets. The distributions are positively skewed and leptokurtic, and the JB test rejects the normality of the log-RVs. This finding is inconsistent with the normality assumption for the error term $u_{t}$ in equation (16), which we retain for tractability.
Remark.
The positive skewness of the log-RVs suggests that more flexible innovations could be considered. In particular, specifying $u_{t}$ as skew-t would allow the model to directly accommodate distributional asymmetry. Incorporating such an extension into the RSV framework would require additional development, so we leave this promising direction for future work.
Table 2: Descriptive statistics of the logarithms of 5-minute RVs for the DJIA and N225.
@@
  • Notes: Standard errors are shown in parentheses. See Table 1 for additional details.